Find an equation of the ellipse that has a center(0,5), a minor axis of length 8, and a vertex at  (0,-1).






1
Expert's answer
2021-02-16T12:38:58-0500

Solution.

Write equation of the ellipse


"\frac{(x-x_0)^2}{a^2}+\frac{(y-y_0)^2}{b^2}=1,"

where "(x_0;y_0)" is a center,

"2a" i "2b" - are axis.

Since the distance from the center of the ellipse to its vertex is 6, then "b=6." The minor axis is 8, so "a=8:2=4."

We will get


"\frac{x^2}{4^2}+\frac{(y-5)^2}{6^2}=1,"


This is the equation of the ellipse. Let's simplify it:


"9x^2+4y^2-40y-44=0."

Answer.

"9x^2+4y^2-40y-44=0."